A star is orbiting a stationary star with an orbital period of 19.0 months. The distance between the two stars is known to be 100 million km. Calculate the mass of the stationary star, assuming a circular orbit.
Answer #1

T =19 months = 19*30*24*60*60 =4.993x10^7 s

R = 100million km =100x106x103 m=1011 m

From keplers law

T2 =(42/GM)R3

M = (42R3/GT2 )

M =[4x3.14x3.14x(1011)3]/[(6.67x10-11)(4.99x107)2]

M =2.4x1029 kg